# RNb-NeuS: Reflectance and Normal-based Multi-View 3D Reconstruction

### v1

@@ Abstract
This paper introduces a versatile paradigm for integrating multi-view reflectance and normal maps acquired through photometric stereo. Our approach employs a pixel-wise joint re-parameterization of reflectance and normal, considering them as a vector of radiances rendered under simulated, varying illumination. This re-parameterization enables the seamless integration of reflectance and normal maps as input data in neural volume rendering-based 3D reconstruction while preserving a single optimization objective. In contrast, recent multi-view photometric stereo (MVPS) methods depend on multiple, potentially conflicting objectives. Despite its apparent simplicity, our proposed approach outperforms state-of-the-art approaches in MVPS benchmarks across F-score, Chamfer distance, and mean angular error metrics. Notably, it significantly improves the detailed 3D reconstruction of areas with high curvature or low visibility.
@@ 1 Introduction
Automatic 3D reconstruction is pivotal in various fields, such as archaeological and cultural heritage for virtual reconstruction, medical imaging for surgical planning and visualization, virtual and augmented reality, games and film production.
@@
In this method, the input reflectance and normal for each pixel are merged into a vector of radiances (or more precisely, luminances, though our approach is extendable beyond the visible spectrum) simulated under arbitrary, varying illumination. In some sense, we project the per-view PS image sets onto a lower-dimensional one which is easier to handle, following well-known methods for robust PS [26]. We then adapt an NVR pipeline to optimize the consistency of these simulations wrt to the scene reflectance and geometry, modeled as the zero-level set of a trained signed distance function (SDF). Coupled with a state-of-the-art PS method such as [8] for obtaining the input reflectance and normals, this approach yields an MVPS pipeline reaching an unprecedented level of fine details, as illustrated in Fig. 1. Besides being the first to exploit reflectance as a prior, our proposed MVPS paradigm is extremely versatile, compatible with any existing or future PS method, whether calibrated or uncalibrated, deep learning-based, or classic optimization procedures.
The rest of this work is organized as follows. State-of-the-art MVPS methods are discussed in Sect. 2. The proposed 3D reconstruction from reflectance and normals is detailed in Sect. 3. Sect. 4 then sketches a proposal for an MVPS algorithm based on this approach. Sect. 5 extensively evaluates this algorithm, before our conclusions are drawn in Sect. 6.
@@ 2 Related work
The first paper to deal with MVPS is by Hernandez et al. [6]. To avoid having to arbitrate the conflicts between the different normal maps, a 3D mesh is iteratively deformed, starting from the visual hull obtained by shape-from-silhouettes, until the recomputed images using the Lambertian model match the original images, while penalizing the discrepancy between the PS normals and those of the 3D mesh. No prior knowledge of camera poses or illumination is required. Under the same assumptions as [6], the method proposed by Park et al. [20, 21] starts from a 3D mesh obtained by SfM (structure-from-motion) and MVS. Simultaneous estimation of reflectance, normals and illumination is achieved by uncalibrated PS, where the normals deduced from the 3D mesh are used to remove the residual ambiguity. In addition, instead of deforming the 3D mesh, the details of the relief are estimated through 2D displacement maps.
@@
Such NeRF-based approaches provide undeniably better 3D reconstructions than classical methods, yet they remain computationally intensive. Recently, Zhao et al. [29] proposed a fast deep learning-based solution to MVPS. For each viewpoint, shading patterns are aggregated so as to create an “artificial shading texture” in textureless regions. The aggregated shading patterns are then matched across viewpoints so that to predict pixel depths and normal maps.
In [12], the authors proposed to complement the solution of [10] by adding a NVR loss term in order to benefit from the reliability of NVR in reconstructing objects with diverse material types. However, this results in a multi-objective optimization comprising three loss terms (besides the Eikonal term). In contrast, we propose a single objective optimization based on an ad hoc re-parametrization which leads to the seamless integration of PS results in standard NVR pipelines. This is detailed in the next paragraph.
@@ 3.2 Reflectance and normal re-parameterization
For the re-reparameterization to be bijective, the reflectance $r_{k}$ must be non-null (a basic assumption in photographic 3D vision), and $\mathsf{L}_{k}$ must be non-singular i.e., the lighting directions must be chosen linearly independent. Then, the original reflectance and normal can be retrieved from the simulated intensities by $r_{k}=\|\mathsf{L}_{k}^{-1}\mathbf{v}(\mathbf{n}_{k},r_{k})\|$ and $\mathbf{n}_{k}=\frac{\mathsf{L}_{k}^{-1}\mathbf{v}(\mathbf{n}_{k},r_{k})}{\|\mathsf{L}_{k}^{-1}\mathbf{v}(\mathbf{n}_{k},r_{k})\|}$. Considering $n>3$ illumination vectors and resorting to the pseudo-inverse operator might induce more robustness, but at the price of losing bijectivity and thus not entirely relying on the PS inputs. We leave this as a possible future work, which might be particularly interesting when the PS inputs are uncertain, or when considering more evolved PBR models involving additional reflectance clues such as roughness, anisotropy or specularity.
In practice, the choice of each arbitrary triplet of light directions $\mathbf{l}_{k,1},\mathbf{l}_{k,2},\mathbf{l}_{k,3}$ can be made so as to minimize the uncertainty on the normal estimate. To this end, the illumination triplet proposed in [4] can be considered. Therein, the authors show that the optimal configuration for three images is vectors that are equally spaced in tilt by $120$ degrees, with a constant slant of $54.74$ degrees (wrt to $\mathbf{n}_{k}$).
@@ 5.1 Experimental setup
We performed our quantitative evaluations using F-score and Chamfer distance (CD), to measure the accuracy of the reconstructed vertices. We also measured the mean angular error (MAE) of the imaged meshes, to evaluate the accuracy of the reconstructed normals wrt the ground truth normals provided in DiLiGenT-MV. We report both the results averaged over all mesh vertices, and those on vertices clustered in two particularly interesting classes, namely high curvature and low visibility areas, as illustrated in Fig. 3. To identify the high curvature areas, we used the library VCGLib [1] and the 3D mesh processing software system Meshlab [3], taking the absolute value of the curvature to merge the convex and concave zones and retaining the vertices whose curvature is higher than $1.6$. To segment the low visibility areas, we summed the boolean visibility of each vertex in each view. Low visibility then corresponds to vertices visible in less than 5 viewpoints, among the 20 ones of DiLiGenT-MV.
@@ 5.3 High curvature and low visibility areas
To highlight the level of details in the 3D reconstructions, Figs. 1 and 6 provide other qualitative comparisons focusing on one small part of each object. Ours is the only method achieving a high fidelity reconstruction on the ear, the knot and the navel of Buddha, and on the spout of Pot2. To quantify this gain, we also report in Table 3 the average CD and MAE over all datasets, yet taking into account only the high curvature and low visibility areas. It is worth noticing that the CD error of PS-NeRF and MVPSNet on high curvature areas increases by $36\%$ and $96\%$, respectively, in comparison with that averaged over the entire set of vertices. Ours, on the contrary, increases by $4\%$ only. Similarly, on low visibility areas their error increases by $78\%$ and $81\%$, and Kaya23 by $46\%$, while ours increases only by $13\%$.
@@ 6 Conclusion
We have introduced a neural volumetric rendering method for 3D surface reconstruction based on reflectance and normal maps, and applied it to multi-view photometric stereo. The proposed method relies on a joint re-parameterization of reflectance and normal as a vector of radiances rendered under simulated, varying illumination. It is very simple to use as it involves a single objective optimization, and it is highly flexible since any photometric stereo method can be used for constructing the input reflectance and normal maps. Therefore, it can be used to integrate results from all existing and future PS methods.
In particular, we have shown empirically that when coupled with a state-of-the-art uncalibrated PS method, our method reaches unprecedented results on the public dataset DiLiGenT-MV in terms of F-score, Chamfer distance and mean angular error metrics. Notably, it provides exceptionally high quality results in areas with high curvature or low visibility.
Its main limitation for now is its computational cost, which we plan to reduce by adapting recent developments within the NeuS2 framework [24].
@@ Acknowledgements.
Baptiste Brument’s doctoral student fellowship is funded by the French Ministry of Higher Education and Research. Robin Bruneau’s doctoral student fellowship is funded by the Danish project PHYLORAMA. This work was partly funded by the French National Research Agency through the LabCom project ALICIA-Vision. Lilian Calvet’s posdoctoral fellowship is supported by the OR-X - a swiss national research infrastructure for translational surgery - and associated funding by the University of Zurich and University Hospital Balgrist.
@@ Appendix A Evaluation
All the state-of-the-art methods were evaluated from the meshes that were kindly provided by their authors. For all evaluated meshes, we eliminated all internal vertices. Then, a mesh upsampling for both estimated and ground truth meshes was then performed in order to achieve a point density of $0.1$ mm. The computations of Chamfer distance and F-score were specifically conducted for distances under $5$ mm in order to mitigate the impact of outliers (inspired by the DTU evaluation [9]).
We observed a few defects in the ground truth meshes from the DiLiGenT-MV dataset in concave areas. Notably, such imperfections are well visible at the back of Bear’s head (Fig. 7) and the spout’s inner area of Pot2 (Fig. 8). Although these areas represent a small amount of vertices, they were discarded in all evaluations so as to avoid penalizing methods which faithfully reconstruct them.
@@ Appendix B Implementation details
We recall that to simulate the radiance values in Step 4 described in Section 4, we choose as lighting triplet the one which is optimal, relatively to the normal $\mathbf{n}_{k}$. Following [4], this optimal triplet is equally spaced in tilt 120 degrees apart with a slant angle of 54.74 degrees. Concretely, the expression of $\mathsf{L}_{k}$ as a function of $\mathbf{n}_{k}$ is written:
@@ Appendix C Additional Results
In this section, we extend the experiments by providing further statistical analysis and qualitative comparisons.
@@
           ———————————————————————————————————————————————— Buddha            ———————————————————————————————————————————————— Reading

### v2

@@ Abstract
This paper introduces a versatile paradigm for integrating multi-view reflectance (optional) and normal maps acquired through photometric stereo. Our approach employs a pixel-wise joint re-parameterization of reflectance and normal, considering them as a vector of radiances rendered under simulated, varying illumination. This re-parameterization enables the seamless integration of reflectance and normal maps as input data in neural volume rendering-based 3D reconstruction while preserving a single optimization objective. In contrast, recent multi-view photometric stereo (MVPS) methods depend on multiple, potentially conflicting objectives. Despite its apparent simplicity, our proposed approach outperforms state-of-the-art approaches in MVPS benchmarks across F-score, Chamfer distance, and mean angular error metrics. Notably, it significantly improves the detailed 3D reconstruction of areas with high curvature or low visibility.
@@ 1 Introduction
Automatic 3D reconstruction is pivotal in various fields, such as archaeological and cultural heritage (virtual reconstruction), medical imaging (surgical planning), virtual and augmented reality, games and film production.
@@
In this method, the input reflectance and normal for each pixel are merged into a vector of radiances simulated under arbitrary, varying illumination. We then adapt an NVR pipeline to optimize the consistency of these simulations wrt to the scene reflectance and geometry, modeled as the zero-level set of a trained signed distance function (SDF). Coupled with a state-of-the-art PS method such as [8] for obtaining the input reflectance and normals, this approach yields an MVPS pipeline reaching an unprecedented level of fine details, as illustrated in Fig. 1. Besides being the first to exploit reflectance as a prior, our proposed MVPS paradigm is extremely versatile, compatible with any existing or future PS method, whether calibrated or uncalibrated, deep learning-based, or classic optimization procedures.
The rest of this work is organized as follows. Sect. 2 discusses state-of-the-art MVPS methods. The proposed 3D reconstruction from reflectance and normals is detailed in Sect. 3. Sect. 4 then sketches a proposal for an MVPS algorithm based on this approach. Sect. 5 extensively evaluates this algorithm, before our conclusions are drawn in Sect. 6.
@@ 2 Related work
The first paper to deal with MVPS is by Hernandez et al. [6]. To avoid having to arbitrate the conflicts between the different normal maps, a 3D mesh is iteratively deformed, starting from the visual hull until the images recomputed using the Lambertian model match the original images, while penalizing the discrepancy between the PS normals and those of the 3D mesh. No prior knowledge of camera poses or illumination is required. Under the same assumptions, Park et al. [20, 21] start from a 3D mesh obtained by SfM (structure-from-motion) and MVS. Simultaneous estimation of reflectance, normals and illumination is achieved by uncalibrated PS, using the normals from the 3D mesh to remove the ambiguity, and estimating the details of the relief through 2D displacement maps.
@@
Such NeRF-based approaches provide undeniably better 3D reconstructions than classical methods, yet they remain computationally intensive. Recently, Zhao et al. [28] proposed a fast deep learning-based solution to MVPS. Aggregated shading patterns are matched across viewpoints so that to predict pixel depths and normal maps.
In [12], the authors proposed to complement [10] by adding a NVR loss term in order to benefit from the reliability of NVR in reconstructing objects with diverse material types. However, this results in a multi-objective optimization comprising three loss terms (besides the Eikonal term). However, similar to [10], the uncertainty-based hyper-parameter tuning does not completely eliminate conflicting objectives, which may induce a loss of fine-scale details. In contrast, we propose a single objective optimization based on an ad hoc re-parametrization which leads to the seamless integration of PS results in standard NVR pipelines. This is detailed in the next paragraph.
@@ 3.2 Reflectance and normal re-parameterization
For the re-reparameterization to be bijective, the reflectance $r_{k}$ must be non-null (a basic assumption in photographic 3D vision), and $\mathsf{L}_{k}$ must be non-singular i.e., the lighting directions must be chosen linearly independent. Then, the original reflectance and normal can be retrieved from the simulated intensities by $r_{k}=\|\mathsf{L}_{k}^{-1}\mathbf{v}(\mathbf{n}_{k},r_{k})\|$ and $\mathbf{n}_{k}=\frac{\mathsf{L}_{k}^{-1}\mathbf{v}(\mathbf{n}_{k},r_{k})}{\|\mathsf{L}_{k}^{-1}\mathbf{v}(\mathbf{n}_{k},r_{k})\|}$. Considering $n>3$ illumination vectors and resorting to the pseudo-inverse operator might induce more robustness but at the price of losing bijectivity and thus not entirely relying on the PS inputs. We leave this as a possible future work, which might be particularly interesting when the PS inputs are uncertain, or when considering more evolved PBR models involving additional reflectance clues such as roughness, anisotropy or specularity.
In practice, the choice of each arbitrary triplet of light directions $\mathbf{l}_{k,1},\mathbf{l}_{k,2},\mathbf{l}_{k,3}$ can be made to minimize the uncertainty on the normal estimate. To this end, the illumination triplet proposed in [4] can be considered. Therein, the authors show that the optimal configuration for three images is vectors that are equally spaced in tilt by $120$ degrees, with a constant slant of $54.74$ degrees (wrt to $\mathbf{n}_{k}$).
@@ 5.1 Experimental setup
We performed our quantitative evaluations using F-score and Chamfer distance (CD), to measure the accuracy of the reconstructed vertices. We also measured the mean angular error (MAE) of the imaged meshes, to evaluate the accuracy of the reconstructed normals wrt the ground truth normals provided in DiLiGenT-MV. We report both the results averaged over all mesh vertices, and those on vertices clustered in two particularly interesting classes, namely high curvature and low visibility areas, as illustrated in Fig. 3.
@@
To identify the high curvature areas, we used the library VCGLib [1] and the 3D mesh processing software system Meshlab [3], taking the absolute value of the curvature to merge the convex and concave zones and retaining the vertices whose curvature is higher than $1.6$. To segment the low visibility areas, we summed the boolean visibility of each vertex in each view. Low visibility then corresponds to vertices visible in less than 5 viewpoints, among the 20 ones of DiLiGenT-MV.
@@ 5.3 High curvature and low visibility areas
To highlight the level of details in the 3D reconstructions, Figs. 1 and 10 provide other qualitative comparisons focusing on one small part of each object. Ours is the only method achieving a high fidelity reconstruction on the ear, the knot and the navel of Buddha, and on the spout of Pot2. To quantify this gain, we also report in Table 3 the average CD and MAE over all datasets, yet taking into account only the high curvature and low visibility areas. It is worth noticing that the CD error of PS-NeRF and MVPSNet on high curvature areas increases by $36\%$ and $96\%$, respectively, in comparison with that averaged over the entire set of vertices. Ours, on the contrary, increases by $4\%$ only. Similarly, on low visibility areas their error increases by $78\%$ and $81\%$, and Kaya23 by $46\%$, while ours increases only by $13\%$.
@@
Bear
Pot2
Input image
Park16
Li19
PS-NeRF
Kaya22
Kaya23
MVPSNet
Ours
GT
Buddha
Cow
Reading
@@ 6 Conclusion
We have introduced a neural volumetric rendering method for 3D surface reconstruction based on reflectance and normal maps, and applied it to multi-view photometric stereo. The proposed method relies on a joint re-parameterization of reflectance and normal as a vector of radiances rendered under simulated, varying illumination. It involves a single objective optimization, and it is highly flexible since any existing or future PS method can be used for constructing the input reflectance and normal maps. Coupled with a state-of-the-art uncalibrated PS method, our method reaches unprecedented results on the public dataset DiLiGenT-MV in terms of F-score, Chamfer distance and mean angular error metrics. Notably, it provides exceptionally high quality results in areas with high curvature or low visibility. Its main limitation for now is its computational cost, which we plan to reduce by adapting recent developments within the NeuS2 framework [24]. Using reflectance uncertainty in addition to that of normal maps offers room for improvement.
@@ Acknowledgements.
This work was supported by the Danish project PHYLORAMA, the ALICIA-Vision project, the IMG project (ANR-20-CE38-0007), the OR-X and associated funding by the University of Zurich and University Hospital Balgrist.
@@ Appendix A Evaluation
All the state-of-the-art methods were evaluated from the meshes that were kindly provided by their authors. For all evaluated meshes, we eliminated all internal vertices. Then, a mesh upsampling for both estimated and ground truth meshes was then performed in order to achieve a point density of $0.1$ mm. The computations of Chamfer distance and F-score were specifically conducted for distances under $5$ mm in order to mitigate the impact of outliers (inspired by the DTU evaluation [9]). We observed a few defects in the ground truth meshes from the DiLiGenT-MV dataset in concave areas. Notably, such imperfections are well visible at the back of Bear’s head (Fig. 7) and the spout’s inner area of Pot2 (Fig. 8). Although these areas represent a small amount of vertices, they were discarded in all evaluations so as to avoid penalizing methods which faithfully reconstruct them.
@@ Appendix B Implementation details
We recall that to simulate the radiance values in Step 4 described in Section 4 of the main paper, we choose as lighting triplet the one which is optimal, relatively to the normal $\mathbf{n}_{k}$. Following [4], this optimal triplet is equally spaced in tilt 120 degrees apart with a slant angle of 54.74 degrees. Concretely, the expression of $\mathsf{L}_{k}$ as a function of $\mathbf{n}_{k}$ is written:
@@ Appendix C Additional Results
In this section, we extend the experiments of the main paper by providing further statistical analysis and qualitative comparisons.
@@
Ours
NeuS
5 views
10 views
15 views
20 views
           ———————————————————————————————————————————————— Buddha Refer to caption
Ours
NeuS
5 views
10 views
15 views
20 views
           ———————————————————————————————————————————————— Reading
@@
Photometric stereo method.
Our method can be employed with any PS method. To illustrate this flexibility, we evaluate the reconstruction accuracy on the Buddha dataset while taking as input the normal maps from CNN-PS [7], used in Kaya22-23 [10, 12], and SDPS-Net [chen2019SDPS_Net], used in PS-NeRF [27], in addition to the one obtained using normal maps from SDM-UniPS [8] reported in the main paper. The results are reported in Table 6. As expected, we observe that the choice of a particular PS technique influences the final outcome, yet our framework consistently improves the results in comparison with previous works, including those based on multi-objective optimizations [10, 12].
CNN-PS SDPS-Net SDM-UniPS Buddha Kaya23 Ours PS-NeRF Ours Ours H. curv. 0.35 0.29 0.51 0.31 0.26 Low curv. 0.24 0.22 0.33 0.25 0.23 All 0.25 0.22 0.34 0.25 0.23
Table 6: Results of our method with different input normals, namely CNN-PS (used in Kaya22-23), SDPS-Net (used in PS-NeRF) and SDM-UniPS. High curvature corresponds to the results averaged over all the vertices whose absolute curvature is higher than $3.3$. Our method shows to perform best irrespective of the PS method being used.
Ablation.
We complete our ablation study with qualitative results on the ear and the knot of Buddha shown in Fig. 10.
W/o reflect.
W/o opt. light.
Ours
GT
Knot
Ear
Figure 10: Qualitative comparison on the knot and the ear of Buddha between our results and those without the use of reflectance and optimal lighting, disabled individually. Our method exhibits better results in both cases.